# Analytical and Numerical Studies of Dark Current in Radiofrequency Structures for Short-Pulse High-Gradient Acceleration

### v2

@@ Acknowledgements.
The authors thank Dr. Haoran Xu and Dr. Omkar Ramachandran for valuable discussions. This research was supported by the U.S. Department of Energy, Office of Science, Office of High Energy Physics under Award DE-SC0021928 and DE-SC0022010.
@@ Appendix A Simplified theory of multipactor in crossed RF fields with closed-form solutions
$\frac{dv_{r}}{dt}=-\frac{e}{m}E_{r}\cos(\omega t+\phi_{0})+\frac{e}{m}v_{0}\sin\alpha_{0}\cdot B_{\theta}\sin(\omega t+\phi_{0})\\ -\left(\frac{e}{m}\right)^{2}B_{\theta}^{2}\,\Delta r\frac{1-\cos(2\omega t+2\phi_{0})}{2}.$ (13)
The oscillatory terms with frequency $2\omega$ are neglected, as their contribution is negligible in the low-field regime. Their influence becomes significant at higher fields, as demonstrated by the phase-space trajectories in Fig. 16. In this example, the approximation remains valid for multipacting orders $N\geq 2$.
@@
$\frac{d^{2}\Delta r}{dt^{2}}+\frac{1}{2}\left(\frac{e}{m}\right)^{2}B_{\theta}^{2}\Delta r=-\frac{e}{m}E_{r}\cos(\omega t+\phi_{0})\\ +\frac{e}{m}v_{0}\sin\alpha_{0}\cdot B_{\theta}\sin(\omega t+\phi_{0}).$ (14)
@@
$\ddot{\Delta r}+\Omega^{2}\Delta r=f(\omega t),$ (15)
@@
This solution holds under the condition $\Omega\ll\omega$, but begins to deviate from full numerical results when $\Omega\gtrsim 0.25\,\omega$. For the system studied here, the condition of $\Omega=\omega$ occurs at $E_{z0}=260.6~\mathrm{MV/m}$, indicating the upper limit of validity for the approximation.
@@
While this simplified approach neglects the radial variation of the transverse fields $E_{r}$ and $B_{\theta}$, and thus cannot predict the secondary electron yield (SEY) with the same accuracy as the full analytical model, it offers a closed-form expression that effectively approximates multipacting resonance conditions, especially in the lower-field regime.
@@
Table 1 summarizes the predicted ranges of $E_{z0}$ corresponding to various multipacting orders, as computed using both the simplified and full analytical models. These ranges are obtained by sweeping over the resonant parameters $z_{\text{res}}$, $\alpha_{\text{res}}$, and $\phi_{\text{res}}$. Notably, good agreement is observed for higher-order multipacting modes ($N\geq 2$), indicating that the simplified model effectively captures the key scaling behavior.

### v3

@@ Acknowledgements.
The authors thank Drs. Chengkun Huang, Omkar Ramachandran and Haoran Xu for valuable discussions. This research was supported by the U.S. Department of Energy, Office of Science, Office of High Energy Physics under Award DE-SC0021928 and DE-SC0022010.
@@ Appendix A Simplified theory of multipactor in crossed RF fields with closed-form solutions
$\frac{dv_{r}}{dt}=-\frac{e}{m}E_{r}\cos(\omega t+\phi_{0})+\frac{e}{m}v_{0}B_{\theta}\sin\alpha_{0}\sin(\omega t+\phi_{0})\\ -\left(\frac{e}{m}\right)^{2}B_{\theta}^{2}\,\Delta r\frac{1-\cos(2\omega t+2\phi_{0})}{2}.$ (14)
The terms oscillating at $2\omega$ are neglected, as their contribution is negligible in the low-field limit. Their influence becomes significant at higher fields, as demonstrated by the phase-space trajectories in Fig. 17. In this example, the approximation is more accurate for multipacting orders $N\geq 2$.
@@
$\frac{d^{2}\Delta r}{dt^{2}}+\frac{1}{2}\left(\frac{e}{m}\right)^{2}B_{\theta}^{2}\Delta r=-\frac{e}{m}E_{r}\cos(\omega t+\phi_{0})\\ +\frac{e}{m}v_{0}B_{\theta}\sin\alpha_{0}\sin(\omega t+\phi_{0}).$ (15)
@@
$\frac{d^{2}\Delta r}{dt^{2}}+\Omega^{2}\Delta r=f(\omega t),$ (16)
@@
This solution holds under the condition $\Omega\ll\omega$, but begins to deviate from full numerical results when $\Omega\gtrsim 0.25\,\omega$. For the system studied here, the condition of $\Omega=\omega$ occurs at $E_{z0}=260.6~\mathrm{MV/m}$, indicating the upper limit of validity for the approximation. Here $E_{z0}$ is computed using parameters obtained by fitting the simulated RF field distribution to Eq. (11).
@@
While this simplified approach neglects the radial variation of the transverse fields $E_{r}$ and $B_{\theta}$, and thus cannot predict the SEY with the same accuracy as the full analytical model, it offers a closed-form expression that effectively approximates multipacting resonance conditions, especially at low fields.
@@
Table 1 summarizes the predicted ranges of $E_{z0}$ corresponding to various multipacting orders, as computed using both the simplified and full analytical models. These ranges are obtained by sweeping over the resonant parameters $z_{\text{res}}$, $\alpha_{\text{res}}$, and $\phi_{\text{res}}$. Notably, good agreement is observed for higher-order multipacting modes ($N\geq 2$), indicating that the simplified theory effectively captures the key scaling behavior.
Figure 18: An example of convergence search for multipacting resonant trajectories with a fixed emission angle $\alpha=42.2^{\circ}$, showing $\Delta r$ versus $z$. The green dot marks the initial release point; the red dot marks the resonant point where the trajectory stabilizes after multiple iterations.
Figure 19: Convergence search over multiple iterations for the example shown in Fig. 18, demonstrating convergence to a resonant, phase-locked trajectory: (a) phase $\phi$, (b) axial position $z$, and (c) SEY $\delta$.
Appendix B Convergence search of multipacting resonant trajectories
As mentioned in Sec. II.3, the analytical theory in this work identifies converged and phase-locked multipacting resonant trajectories, in contrast to previous studies [38]. In this appendix, we present an example of the convergence search process, shown in Fig. 18. In this example, we fix the emission angle $\alpha$ while the other variables are allowed to vary. Each iteration integrates one trip from emission to impact, while the RF phase $\phi$ and axial position $z$ at emission are set to the previous impact values. The earlier paths are not resonant, while the trajectory ultimately converges to a phase-locked stable orbit. These successive iterations do not represent the physical time evolution of secondary electron motion over multiple impacts; rather, they constitute a numerical procedure to locate phase-locked resonance conditions. Convergence is declared when, for ten consecutive iterations, the changes in $\phi$, $z$, and SEY $\delta$ between iteration $k$ and $k+1$ each fall below $10^{-6}$, while $\delta$ remains greater than unity:
${\color[rgb]{0,0,0}|\phi_{k+1}-\phi_{k}|<10^{-6},\quad|z_{k+1}-z_{k}|<10^{-6},}\\ {\color[rgb]{0,0,0}|\delta_{k+1}-\delta_{k}|<10^{-6},\quad\delta_{k}>1.}$
Figure 19 illustrates the convergence of these quantities, confirming the establishment of the resonance condition.